Feet-First Industries plans to sell 7,750 sleds at $85 each in the coming year. Variable cost is 60 percent of the sales price. Fixed factory overhead equals $52,940 and fixed selling and administrative expense equals $34,780.

a. Calculate the units that Feet-First must sell in order to break even.

Answer

b. Calculate the sales revenue that Feet-First must earn to break even by using the contribution margin.

$Answer

c. Confirm your answer in requirement b, by muliplying the number of break-even units in requirement a by the unit sales price.

$Answer
Answer #1

a) Break even = Fixed cost/Contribution margin per unit = (52940+34780)/34 = 2580

b) Contribution margin ratio =40%

Break even sales = 87720/.40 = 219300

c) Break even sales = 2580*85 = 219300